# Kekulé order from diffuse nesting near higher-order Van Hove points

### v1

@@ B Quasiparticle bandstructures
From the GL analysis, the importance of the global phase of the order parameter becomes directly apparent. In the GL functional, this is reflected by the non-vanishing third order terms due to the relation $3K=\Gamma$. For $\phi_{1}\in\{1/3\pi,\pi,5/3\pi\}$ we have two stronger bonded and one weaker bonded triangle in the $\sqrt{3}\times\sqrt{3}$ unit cell, while for $\phi_{2}\in\{0,2\pi/3,4\pi/3\}$ the situation is reversed (cf. insets of Fig. SM 9). The structure of the leading FRG instability locks the $\pi$ phase shift between the onsite and bond components (cf. Fig. SM 7). This has direct consequences for the quasiparticle bandstructures obtained by diagonalizing the MF Hamiltonian in Eq. SM 5, where the quasiparticle gap at the $\Gamma$ point (formerly $K^{(\prime)}$) is significantly enhanced for $\phi_{1}$. The Kekulé charge order destroys the higher order nature of the vHs with a flat dispersion along the $M=M_{r}-K_{r}$ direction and induces a transition to an ordinary vHs. As a result, significant spectral weight is shifted away from the Fermi level, and the system lowers its free energy considerably.
Figure SM 9: Quasiparticle bandstructure of the breathing kagome for MF states with different global phases $\phi$. Grey lines indicate the bandstructure of the breathing kagome model backfolded from the original BZ (grey) to the reduced BZ (black) as shown in the inset. Black lines show the quasiparticle bandstructure for the order parameter in Eq. SM 11 with $A=0.1t$ and $\phi=\pi(0)$ in the left (right) panel. The insets display the order parameter in real space according to the convention in Fig. SM 7 with the original (enlarged) unit cell indicated in grey (black). Red (cyan) corresponds to negative (positive) values of the order parameter on the respective sites and bonds.
@@
It is curious to draw attention to the real space similarities between our charge ordered kagome Kekulé phase and those previously studied in honeycomb models. The kagome lattice can be constructed as the line graph of a honeycomb lattice. Inspecting the line graph reconstruction of the onsite part of the obtained $K$-order reveals a Kekulé-Y order [136] on the effective honeycomb structure, as shown in Fig. SM 10.
Figure SM 10: Line graph reconstruction of the kagome lattice revealing a honeycomb structure. The bonds are colored corresponding to the onsite values of the associated onsite kagome Kekulé order forming a Kekulé-Y order on the effective honeycomb lattice.
@@
We analyze the quantum metric for different magnitudes of breathing, corresponding to varying gap sizes between the band hosting the HOVHS and the two residual bands. A linear interpolation of the kinetic parameters is employed between the non-breathing HOVHS kagome lattice [Fig. SM 2(b)] and the breathing case [Fig. SM 2(a)]. The Brillouin zone integral over the quantum metric $g(\boldsymbol{k})$ is given by
$\Omega_{\mathrm{I}}=\int_{\mathrm{BZ}}\frac{d^{2}\boldsymbol{k}}{V_{\mathrm{BZ}}}\,\mathrm{Tr}\,[g(\boldsymbol{k})].$ (SM 15)
The quantity $\Omega_{\mathrm{I}}$ is related to the gauge independent spread of Wannier functions [119, 120].
@@
Figure SM 11: Quantum metric (insets) and Brillouin zone integral over the quantum metric for linear interpolation between non-breathing Fig. SM 2(b) and breathing Fig. SM 2(a) kagome lattice.
@@
Fig. SM 11 displays the evolution of $\Omega_{\mathrm{I}}$ from an almost non-breathing configuration ($t_{1}/t_{2}\approx 1$) to the breathing model discussed in the main text. The decrease in $\Omega_{\mathrm{I}}$ indicates a stronger localization of the Wannier functions. The insets show the quantum metric in the BZ for $t_{1}/t_{2}\approx 1.06$ and $t_{1}/t_{2}\approx 2.24$. While being highly localized in momentum space for the nearly non-breathing case, the quantum metric becomes lower in magnitude and more spread out across the BZ in the breathing scenario, corroborating the behavior observed in $\Omega_{\mathrm{I}}$.

### v2

@@ B Quasiparticle bandstructures
From the GL analysis, the importance of the global phase of the order parameter becomes directly apparent. In the GL functional, this is reflected by the non-vanishing third order terms due to the relation $3K=\Gamma$. For $\phi_{1}\in\{1/3\pi,\pi,5/3\pi\}$ we have two stronger bonded and one weaker bonded triangle in the $\sqrt{3}\times\sqrt{3}$ unit cell, while for $\phi_{2}\in\{0,2\pi/3,4\pi/3\}$ the situation is reversed (cf. insets of Fig. SM 8). The structure of the leading FRG instability locks the $\pi$ phase shift between the onsite and bond components (cf. Fig. SM 6). This has direct consequences for the quasiparticle bandstructures obtained by diagonalizing the MF Hamiltonian in Eq. SM 6, where the quasiparticle gap at the $\Gamma$ point (formerly $K^{(\prime)}$) is significantly enhanced for $\phi_{1}$. The Kekulé charge order destroys the higher order nature of the vHs with a flat dispersion along the $M=M_{r}-K_{r}$ direction and induces a transition to an ordinary vHs. As a result, significant spectral weight is shifted away from the Fermi level, and the system lowers its free energy considerably.
Figure SM 8: Quasiparticle bandstructure of the breathing kagome for MF states with different global phases $\phi$. Grey lines indicate the bandstructure of the breathing kagome model backfolded from the original BZ (grey) to the reduced BZ (black) as shown in the inset. Black lines show the quasiparticle bandstructure for the order parameter in Eq. SM 12 with $A=0.1t$ and $\phi=\pi(0)$ in the left (right) panel. The insets display the order parameter in real space according to the convention in Fig. SM 6 with the original (enlarged) unit cell indicated in grey (black). Red (cyan) corresponds to negative (positive) values of the order parameter on the respective sites and bonds.
@@
Our $\sqrt{3}\times\sqrt{3}$ order exhibits several defining features of Kekulé order, such as translation symmetry breaking, tripling of the unit cell, gapping of the Dirac point, momentum-space structure, and bond-order modulation, which go beyond a generic $\sqrt{3}\times\sqrt{3}$ order.
While a protecting symmetry to either site or bond ordering is absent in our model, and any interpolation between site and bond contribution can be found in a given charge or spin density wave susceptibility channel, we find the bond ordering part to largely dominante within our $\sqrt{3}\times\sqrt{3}$ order.
Since the concept of Kekulé order is etymologically based on the properties of the honeycomb lattice, it is curious to draw attention to the real space similarities between our charge ordered kagome Kekulé phase and those previously studied in honeycomb models. The kagome lattice can be constructed as the line graph of a honeycomb lattice. Inspecting the line graph reconstruction of the onsite part of the obtained $K$-order reveals a Kekulé-Y order [28] on the effective honeycomb structure, as shown in Fig. SM 9. From these shared properties, we conclude that the observed order can be rigorously identified as Kekulé order.
Figure SM 9: Line graph reconstruction of the kagome lattice revealing a honeycomb structure. The bonds are colored corresponding to the onsite values of the associated onsite kagome Kekulé order forming a Kekulé-Y order on the effective honeycomb lattice.
@@
A Quantum metric
The quantum metric $g_{\mu\nu}^{(n)}(\boldsymbol{k})$ for band $n$ is defined as the real part of the quantum geometric tensor $Q_{\mu\nu}^{(n)}(\bm{k})=\langle\partial_{k_{\mu}}u_{n}(\bm{k})|\left(1-|u_{n}(\bm{k})\rangle\langle u_{n}(\bm{k})|\right)|\partial_{k_{\nu}}u_{n}(\bm{k})\rangle$ with the Bloch function of the $n$-th band $|u_{n}(\bm{k})\rangle$. It defines the infinitesimal distance between nearby states in Hilbert space as $ds_{n}^{2}(\bm{k})=g_{\mu\nu}^{(n)}(\bm{k})\,dk_{\mu}\,dk_{\nu}$. Its Brillouin zone integral
$\Omega_{\mathrm{I}}=\int_{\mathrm{BZ}}\frac{d^{2}\boldsymbol{k}}{V_{\mathrm{BZ}}}\,\mathrm{Tr}\,[g(\boldsymbol{k})]$ (SM 16)
provides a lower bound for the spread of Wannier functions [134, 121]. For our purposes, this is important because an isolated band with a small value of $\Omega_{\mathrm{I}}$ admits a description in terms of well-localized Wannier functions. Consequently, the system can be accurately reduced to a single-orbital tight-binding model for that band, which we require to construct a meaningful low-energy effective model (Section C).
@@ A Quantum metric
We analyze $\Omega_{\mathrm{I}}$ for different magnitudes of breathing, corresponding to varying gap sizes between the band hosting the HOVHS and the two residual bands. A linear interpolation of the kinetic parameters is employed between the non-breathing HOVHS kagome lattice [Fig. SM 2(b)] and the breathing case [Fig. SM 2(a)].
@@
Figure SM 10: Quantum metric (insets) and Brillouin zone integral over the quantum metric for linear interpolation between non-breathing Fig. SM 2(b) and breathing Fig. SM 2(a) kagome lattice.
@@
Fig. SM 10 displays the evolution of $\Omega_{\mathrm{I}}$ from an almost non-breathing configuration ($t_{1}/t_{2}\approx 1$) to the breathing model discussed in the main text. The decrease in $\Omega_{\mathrm{I}}$ indicates a stronger localization of the Wannier functions. The insets show the quantum metric in the BZ for $t_{1}/t_{2}\approx 1.06$ and $t_{1}/t_{2}\approx 2.24$. While being highly localized in momentum space for the nearly non-breathing case, the quantum metric becomes lower in magnitude and more spread out across the BZ in the breathing scenario, corroborating the behavior observed in $\Omega_{\mathrm{I}}$.
B Effects of sublattice polarization
The lattice geometry of the unmodified kagome model leads to a complete (partial) sublattice polarization at the pure $p$- (mixed $m$-) type VHS at the $M$ points, protected by $\sigma_{v}$ mirror symmetry. Upon opening an infinitesimal gap, which is required to define a lower band and thus a quantum metric for it, the lower bound $\Omega_{\mathrm{I}}$ of the Wannier function spread is evaluated to be large. This indicates that a single-orbital description of the model is not possible.
In the breathing configuration, the $\sigma_{v}$ symmetry is explicitly broken on the single-particle level and the FS states feature support on all three sublattices throughout the Fermi surface. Hence, the nesting analysis in Fig. SM 4 does not require a detailed examination of the electronic eigenstates on the FS. As the breathing introduces a trivial gap in the spectrum which is accompanied by decreased $\Omega_{\mathrm{I}}$ (see Fig. SM 10), it permits a single-orbital description of the lowest band hosting the HOVHS. This naturally excludes any sublattice polarization effects from emerging in the low-energy description of the system.
Figure SM 11: Fermi surface nesting in the non-breathing scenario. (a) C sublattice weight along the Fermi surface, indicated by color. The sublattice polarization varies smoothly along the FS and is hence present over an extended region around the $M$-point, where a vanishing C weight is enforced by $C_{2}$ symmetry. (b) Original (red) and $K$-shifted (blue) Fermi surface. The opacity indicates again contribution of sublattice $C$ to the FS states. The FS regions with most weight on C only nests with FS segments with little C contribution (indicated by the black rectangle) reproducing the well-known sublattice interference effect also for $K$-nesting. (c) Same as (b) but now with different sublattice contribution on the shifted Fermi surface, indicating the available scattering processes for nearest-neighbor interactions. Now FS states with high occupation on both the original and shifted FS nest to drive a $K$-ordering.
In the fully $C_{6v}$ symmetric non-breathing HOVHS scenario, the HOVHS is of mixed sublattice type and the FS around the three inequivalent $M$-points becomes depopulated of weight from a specific sublattice (cf. Fig. SM 11(a)). As in the unmodified kagome model, this is accompanied by an increased lower bound of the Wannier function spread (see Fig. SM 10). It is worth noting that due to the continuous evolution of the FS states, especially in the FS segments with large DOS close to the HOVHS points along the $\sigma_{v}$ mirror plane, sublattice polarization is strong even away from the singularity. This necessitates a sublattice-resolved analysis of the $\boldsymbol{q}=K$ nesting that differs substantially between the identical FS shapes in Fig. SM 4: The overlap region highlighted in Fig. SM 11(b) indicates a dominant FS segment of low sublattice weight, which is nested with regions of high weight of the same sublattice. These states are therefore inaccessible to local scattering events and require a mediation by a NN repulsion that mixes sublattices (compare Fig. SM 11(b) and (c)). This scenario for the $K$-nesting is exactly analogous to the well-known sublattice interference effect for the $M$ point VHS nesting in the kagome Hubbard model [56]. Consequently, we inspect a suppression of critical scales and a disappearing $K$-phase in the small $V_{12}$ regime in Fig. SM 3(c). For large NN repulsion, we recover the usual SC order known from FRG studies on the conventional kagome Hubbard model [117, 55, 91]. While sublattice interference, as originally proposed, favors bond order formation by enhancing longer-range correlations, the simultaneous suppression of the critical scale disfavors the Kekulé bond order, which we have shown to arise from diffuse nesting.